A quantity of gas occupies a volume of 548 mL at a temperature of 27 degrees C. What temperature (in Kelvin) will the volume of the gas be at 600 mL, assuming no change in pressure?


1
Expert's answer
2021-05-26T07:07:53-0400

"\frac {V_1} {T_1} =\frac {V_2}{T_2}"


"T_2=\frac {300×600}{548}"


"T_2=328.5K"